The least-squares regression Ime for predicting the nbeye pnce rom ue con ph y 4.8904 + 2.8678x. Predict the ribeye price in a month when the corn price was $6.40 per bushel. A)$12.79 per Ib B)$15.30 per lb 9$13.46 per lb D) $18.35 10) For the following data set, compute the coefficient of determination. x 3 7 1 4 6 2 V9 20 21 19 22 12 A) 0.835 B) 0.594 C) 0.165 D) 0.406 11) For the following data set, how much of the variation in the outcome variable is exp least-squares regression line? y 24 17 10 18 20 23 A)31,1% B) 90.3% C) 68.9% D) 9.7% Pationt: Mr. Colby Smith

#10 and #11
Answer #1

Ans:

10)

x y xy x^2 y^2
1 3 9 27 9 81
2 7 20 140 49 400
3 1 21 21 1 441
4 4 19 76 16 361
5 6 22 132 36 484
6 2 12 24 4 144
Total 23 103 420 115 1911

Correlation coefficient,r=(6*420-23*103)/sqrt((6*115-23^2)*(6*1911-103^2))=0.4065

Coefficient of determination,r^2=0.4065^2=0.165

Option C is correct.

11)

x y xy x^2 y^2
1 3 24 72 9 576
2 7 17 119 49 289
3 1 10 10 1 100
4 5 18 90 25 324
5 8 20 160 64 400
6 4 23 92 16 529
Total 28 112 543 164 2218

Correlation coefficient,r=(6*543-28*112)/SQRT((6*164-28^2)*(6*2218-112^2))=0.3121

Coefficient of determination,r^2=0.3121^2=0.097

So,correct option is D (9.7%)